OpenStudy (kc_kennylau):

How to measure the hydronium ion concentration of water? :) EDIT: like I know it's 1x10^-7M but how

OpenStudy (kc_kennylau):

@ganeshie8 @myininaya @.Sam. @thomaster @timo86m @dan815 @Compassionate @shamil98 @TheViper @FutureMathProfessor @kewlgeek555 @AllTehMaffs

OpenStudy (kc_kennylau):

@Vincent-Lyon.Fr Aide-moi si tu peux :)

OpenStudy (anonymous):

\[ 2 H_2O \overset{\leftarrow}{\rightarrow} OH- + H_3O^+\] In pure water there are an equal number of hydroxide ions as hydronium ions, and pH = 7 pOH = 14 - 7 = pOH =7 means \[7=-\log{a_{OH-}}\] and you're looking for aOH

OpenStudy (vincent-lyon.fr):

Measure its pH. If you find pH = 7, then it means the concentration is \(10^{-7}\) mol/L

OpenStudy (anonymous):

Top reaction is auto-dissociation in water

OpenStudy (kc_kennylau):

But pH comes from the concentration, you can't reverse it...

OpenStudy (anonymous):

?

OpenStudy (kc_kennylau):

like you know water's pH is 7 because its hydonium ion concentration is 1x10^-7

OpenStudy (anonymous):

pH hydrogen ion concentration, and pOH s hydronium ion concentration - in water they're related by pH + pOH = 14 I don't understand what's being reversed however...

OpenStudy (kc_kennylau):

how do you know that pH+pOH=14...

OpenStudy (anonymous):

It's from an equilibrium constant that governs the above chemical equation (I'm not very well versed in it - I was just always taught that was true) http://en.wikipedia.org/wiki/PH#pOH from \[pH = pK_w - pOH\] and \[pK_w \approx 14 \ \text{at room(ish) temperature}\]

OpenStudy (kc_kennylau):

ok...

OpenStudy (aaronq):

with a pH meter?

OpenStudy (anonymous):

I appear to have been incorrect, I'm very sorry - pH seems to effectively be a measure of the concentration of hydronium ions, so since the pH of water is 7, then the concentration of hydronium ions is \[ [H_3O^+] = 10^{-pH}\] which was the answer you were looking for. Again, apologies.

OpenStudy (anonymous):

Also, above, when I mentioned that pOH was hydronium I mistyped - pOH is hydroxide ions, (OH- ions). That's why everything I said was so confusing I think (also because it was mostly incorrect. So again, apologies).

OpenStudy (kc_kennylau):

but I mean how do you know that water's pH is 7 in the first place

OpenStudy (anonymous):

I'm not entirely sure how to mathematically derive it - it's just generally what pure water is taken to be. Sorry that it's a cop out answer :P

OpenStudy (kc_kennylau):

so the pH of water is defined to be 7?

OpenStudy (aaronq):

yes it is, and (in a more rigorous approach) it's based on the activity of hydronium ions, not their concentration.

OpenStudy (kc_kennylau):

ic